The sides of a right angle triangle PQR are PQ = 7cm, QR = 25cm angle P = 90° respectively then what is tanQ – tanR.
 

Grade:12

4 Answers

SAIMANIKANTA
8 years ago
PQ = 7cm and QR = 25 cm.
PQR is a right angle triangleBy using Pythagoras theorem
PR2 =QR2 - PQ2
PR2=576
PR=24
Tan Q=PR/PQ=24/7
Tan R=PQ/PR=7/24
Tan Q-Tan R=527/168
 
SAI SANDY
8 years ago
 
PQ = 7cm and QR = 25 cm.
PQR is a right angle triangleBy using Pythagoras theorem
PR2 =QR2 - PQ2
PR2=576
PR=24
Tan Q=PR/PQ=24/7
Tan R=PQ/PR=7/24
Tan Q-Tan R=527/168 .
Tinku
2 years ago
PQ = 7cm and QR = 25 cm.
PQR is a right angle triangleBy using Pythagoras theorem
PR2 =QR2 - PQ2
PR2=576
PR=24
Tan Q=PR/PQ=24/7
Tan R=PQ/PR=7/24
Tan Q-Tan R=527/168 .
 
Vikas Amritiya
2 years ago
PQ = 7cm and QR = 25 cm.
PQR is a right angle triangleBy using Pythagoras theorem
PR2 =QR2 - PQ2
PR2=576
PR=24
Tan Q=PR/PQ=24/7
Tan R=PQ/PR=7/24
Tan Q-Tan R=527/168